# Refinement-based Christoffel sampling for least squares approximation in non-orthogonal bases

### v1

@@ Abstract
We introduce a refinement-based Christoffel sampling (RCS) algorithm for least squares approximation in the span of a given, generally non-orthogonal set of functions $\Phi_{n}=\{\phi_{1},\dots,\phi_{n}\}$. A standard sampling strategy for this problem is Christoffel sampling, which achieves near-best approximations in probability using only $\mathcal{O}(n\log(n))$ samples. However, it requires i.i.d. sampling from a distribution whose density is proportional to the inverse Christoffel function $k_{n}$, the computation of which requires an orthonormal basis. As a result, existing approaches for non-orthogonal bases $\Phi_{n}$ typically rely on costly discrete orthogonalization. We propose a new iterative algorithm, inspired by recent advances in approximate leverage score sampling, that avoids this bottleneck. Crucially, while the computational cost of discrete orthogonalization grows proportionally with $\|k_{n}\|_{L^{\infty}(X)}$, the cost of our approach increases only logarithmically in $\|k_{n}\|_{L^{\infty}(X)}$. In addition, we account for finite-precision effects by considering a numerical variant of the Christoffel function, ensuring that the algorithm relies only on computable quantities. Alongside a convergence proof, we present extensive numerical experiments demonstrating the efficiency and robustness of the proposed method.
@@ 1 Introduction
We consider the problem of approximating a function $f\in L^{2}_{\rho}(X)$, defined on $X\subseteq\mathbb{R}^{d}$, using weighted least squares fitting in the span of a given basis $\Phi_{n}=\{\phi_{1},\dots,\phi_{n}\}$
@@
where the set of sample points $\{x_{k}\}_{k=1}^{m}$ and the weights $\{w_{k}\}_{k=1}^{m}$ can be chosen freely. The objective is to attain near-best approximation errors while using a small number of sample points. For simplicity of exposition, we assume that the functions $\phi_{i}$ are linearly independent so that the dimension of $\operatorname{span}(\Phi_{n})$ equals $n$; however, the discussed methods also apply to linearly dependent sets of functions.
A common strategy is to use sample points that are drawn i.i.d. from some sampling distribution, allowing for an analysis using tools from random matrix theory. This led to the sampling distribution $\mu_{n}$, first proposed for (total degree) polynomial spaces [21] and later generalized [9], which achieves a near-optimal approximation error in probability using only $m=\mathcal{O}(n\log n)$ samples. This is close to the optimal case of interpolation, which corresponds to $m=n$. The sampling distribution $\mu_{n}$ is defined by
@@
is known as the inverse of the Christoffel function [31] and, hence, this sampling strategy is often referred to as Christoffel sampling [1].
@@
In earlier work [15, 2, 30], this issue was addressed by discretely orthogonalizing the basis $\Phi_{n}$ on a dense grid. More specifically, we can equivalently characterize the inverse Christoffel function $k_{n}$ using the non-orthogonal basis $\Phi_{n}$ by
@@
where the points $\{t_{k}\}_{k=1}^{\ell}$ are sampled i.i.d. from the measure $\rho$. Following [15, 1], one needs $\ell=\mathcal{O}(\|k_{n}\|_{L^{\infty}(X)}\log(n))$ for $G_{\ell}$ to approximate $G$ well with high probability. Although this is typically an offline cost—computed once for every $\Phi_{n}$ and independently of the function to be approximated—it can still be prohibitively expensive, especially when the Christoffel function $k_{n}$ is sharply peaked. Notably, this is precisely the regime where Christoffel sampling offers significant advantages over standard least squares fitting using samples drawn from $\rho$. When $\ell$ is too small, the efficiency of the resulting Christoffel sampling procedure is limited.
We adopt a different strategy inspired by [10]; Section 1.2 explains how this discrete method relates to our continuous setting. We aim to construct a function $u$ that satisfies $u(x)\geq k_{n}(x),\forall x\in X$ while $\|u\|_{L^{1}_{\rho}(X)}\leq Cn$ for some small constant $C\geq 1$. When such a function is known, it suffices to draw $m=\mathcal{O}(\|u\|_{L^{1}_{\rho}(X)}\log(n))=\mathcal{O}(n\log(n))$ samples from the sampling distribution $\mu_{u}$ defined by
@@
in order to obtain near-optimal weighted least squares approximations in expectation. We construct such a function $u$ efficiently using an iterative method in which, at each step, we sample points from the current $u$ and thereafter use these samples to update and improve $u$.
In particular, we show that $\|u\|_{L^{1}_{\rho}(X)}$ decreases by a constant factor in each iteration until it is close to $\|k_{n}\|_{L^{1}_{\rho}(X)}=n$, using only $\mathcal{O}(n\log(n))$ samples per step. Starting from the initialization $u(x)=\|k_{n}\|_{L^{\infty}(X)}$, it follows that the total number of samples required to obtain an adequate $u$ is of the order $\mathcal{O}(\log(\|k_{n}\|_{L^{\infty}(X)})n\log(n))$. Therefore, in contrast to the method described above, the computational complexity of this approach depends only logarithmically on $\|k_{n}\|_{L^{\infty}(X)}$, rather than linearly. Inspired by [10, Algorithm 3], the algorithm is termed refinement-based Christoffel sampling (RCS).
@@ 1.1 Effects of finite-precision arithmetic
Another issue related to least squares fitting with non-orthogonal bases is the influence of numerical rounding errors. Even if a basis is linearly independent from an analytical perspective, it can be indistinguishable from a linearly dependent set in finite-precision arithmetic. Numerically computed least squares fits using such a basis behave either explicitly or implicitly as $\ell^{2}$-regularized approximations with a regularization parameter $\epsilon\propto\epsilon_{\text{mach}}\sqrt{\|G\|_{2}}$ due to rounding errors [4, 5, 23]. Whereas this numerical regularization has a negligible influence if the basis is (nearly) orthogonal, it becomes significant when $\kappa(G)\geq 1/\epsilon_{\text{mach}}^{2}$, in which case $\Phi_{n}$ is called numerically redundant [23].
Using similar tools from random matrix theory, it was found in [23] that the sampling distribution $\mu_{n}^{\epsilon}$ defined by
@@ 1.2 Connection to the discrete setting
Christoffel sampling is widely recognized as being closely related to leverage score sampling, a well-established technique in numerical linear algebra [18, 1]. Given an $m\times n$ matrix $A$ with $m>n$, the $j$-th statistical leverage score $\ell_{j}$ is defined by
@@
where $U$ denotes the $m\times n$ matrix containing the left singular vectors of the matrix $A$ [16]. Leverage scores identify which rows of a matrix are most important and are typically used to subsample a tall-and-skinny matrix $A$ before solving a least squares problem $Ax\approx b$. Other applications include, but are not limited to, low-rank approximation, matrix completion and kernel-based machine learning [17].
@@
The numerical Christoffel function $k_{n}^{\epsilon}$ can similarly be seen as a continuous analogue of ridge leverage scores [23], which arise in the context of $\ell^{2}$-regularized least squares problems [6].
Challenges similar to those encountered in Christoffel sampling also appear in the context of leverage scores. In particular, computing leverage scores requires the full singular value decomposition of $A$, which is as computationally expensive as solving the original least squares problem. Consequently, there has been substantial research on fast computation—or rather, approximation—of leverage scores. One approach introduced in [10, 11] consists of computing a tight upper bound to the (rigde) leverage scores via an efficient iterative scheme. This approach can be extended to the continuous setting and forms the foundation of this paper.
@@ 1.3 Other related work
Other methods have been proposed for efficiently computing the Christoffel function from a non-orthogonal basis. For example, in  [15, §4] a multilevel strategy is introduced that approximates the inverse Christoffel function through successive refinements based on nested subspaces. However, this approach relies on rather strong assumptions about these subspaces and does not account for the effects of finite-precision arithmetic. Another iterative method is presented in [33], which requires the strong assumption $\lambda_{\min}(G)\gg\epsilon_{\text{mach}}$—where $\lambda_{\min}$ indicates the smallest nonzero eigenvalue—for computability.
Moreover, we note that optimizations have been studied to improve upon the Christoffel sampling scheme. These strategies reduce the number of required samples from $\mathcal{O}(n\log(n))$ to $\mathcal{O}(n)$, typically by leaving the i.i.d. setting. For instance, in [7] a subsampling scheme is introduced that is able to compress a good sampling set, while the algorithm in [14] is designed to directly sample from a distribution that incorporates both the Christoffel function and effective resistance. Implementing the latter can be done via rejection sampling with draws from the Christoffel function, as mentioned in [14, §5].
The link between leverage scores and the Christoffel function is not new; a more extensive comparison can be found in [29, §2]. Furthermore, in the discrete setting many more algorithms have been proposed to efficiently approximate leverage scores, such as [16] which uses random sketching to speed up the computations. Viewing the object $\begin{bmatrix}\phi_{1}&\phi_{2}&\dots&\phi_{n}\end{bmatrix}$ as the limit of a tall-and-skinny matrix was popularized in the Matlab package chebfun, where it is referred to as a quasimatrix [19].
@@ 1.4 Overview of the paper
The remainder of the paper is organized as follows. In Section 2, the symbols and notation used throughout the paper are summarized. In Section 3, we introduce the proposed iterative algorithm, outline its main objective and provide an initial discussion of its computational performance. Section 4 presents a theoretical analysis of the method, including a proof of its convergence properties. Finally, Section 5 is devoted to extensive numerical experiments that illustrate the efficiency and robustness of our approach.
@@ 2 Symbols and notation
in (1). The set $\Phi_{n}=\{\phi_{1},\dots,\phi_{n}\}\subset L^{2}_{\rho}(X)$ is a given set of (non-orthogonal) basis functions. The vector of basis evaluations at a point $x\in X$ is denoted by $\phi(x)=\begin{bmatrix}\phi_{1}(x)&\dots&\phi_{n}(x)\end{bmatrix}^{\top}\in\mathbb{C}^{n}$. The continuous Gram matrix associated with $\Phi_{n}$ is an $n\times n$ matrix $G$ defined by
@@ 3.1 Main objective
The goal is to identify a small number of samples $\{x_{k}\}_{k=1}^{m}$ and weights $\{w_{k}\}_{k=1}^{m}$ so that the discrete least squares fit to a function $f$ in a given basis $\Phi_{n}$ has near-optimal approximation error. Our approach to obtain such samples is to compute a function $u$ that satisfies
@@
closely approximates the continuous Gram matrix $G$ using only $m=\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))$ samples. This is formalized in Lemma 3.1, which generalizes [23, Theorem 10].
@@
Let $\Phi_{n}=\{\phi_{i}\}_{i=1}^{n}\subset L^{2}_{\rho}(X)$ be a basis, where $X\subseteq\mathbb{R}^{d}$, $\epsilon\geq 0$, and $u$ be an integrable function satisfying
@@
This follows from a straightforward generalization of the proof of [23, Theorem 10] using a generic $\Delta$ instead of $\Delta=1/2$, and from
@@
Throughout this paper we will assume that $k_{n}^{\epsilon}(x)>0,\forall x\in X$. This is a standard assumption in Christoffel sampling schemes and is satisfied if at every point $x\in X$ at least one basis function is nonzero. This condition is mild and, for example, holds whenever the constant function is included in the basis. An alternative is to slightly modify the Christoffel density, as explained in [1, §6.1].
Using [23, Theorem 2], this immediately implies that numerically computed least squares fits are stable and accurate.
@@
where the term $\epsilon^{2}/\lambda_{1}(G)$ is negligble when $\Phi_{n}$ is close to orthogonal. From here, it follows that the discrete least squares error is near-optimal using $\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))=\mathcal{O}(n\log(n))$ samples, by arguments analogous to those used in standard results such as [9, Theorem 2.1].
@@ 3.2 Iterative scheme
The core strategy for constructing a function $u$ that satisfies (4) is an adaptation of the (ridge) leverage score sampling algorithm introduced in [10, 11] to the continuous setting. Suppose we have a candidate function $u$ that upper bounds $k_{n}^{\epsilon}$ but with $\|u\|_{L^{1}_{\rho}(X)}\gg n^{\epsilon}$. Drawing fewer than $\mathcal{O}(\|u\|_{L^{1}_{\rho}(X)}\log(n^{\epsilon}))$ samples from the associated measure $\mu_{u}$ (3) is generally insufficient for accurate least squares fitting. However, these samples can still be used to improve uu—that is, to construct a new function $u^{(new)}$ that still upper bounds $k_{n}^{\epsilon}$ but has a reduced $L_{\rho}^{1}(X)$-norm. This procedure is applied iteratively, producing successively tighter upper bounds to $k_{n}^{\epsilon}$ until the requirement (4) is met.
@@
with high probability. Thus, for given $C_{1}$ and $\Delta$, choosing $C_{2}$ appropriately ensures a geometric decrease in $L_{\rho}^{1}(X)$-norm of $u$, using only $\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))$ samples in each iteration. Note that we don’t have to compute $m$ or $\alpha$; we simply draw $\lceil\alpha m+1\rceil$ samples in each iteration.
@@
$\Phi_{n}\subset L^{2}_{\rho}(X)$, (upper bounds on) $\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}$ and $n^{\epsilon}$

### v2

@@ Abstract
We introduce a refinement-based Christoffel sampling (RCS) algorithm for least squares approximation in the span of a given, generally non-orthogonal set of functions $\Phi_{n}=\{\phi_{1},\dots,\phi_{n}\}$. A standard sampling strategy for this problem is Christoffel sampling, which achieves near-best approximations in probability using only $\mathcal{O}(n\log(n))$ samples. However, it requires i.i.d. sampling from a distribution whose density is proportional to the inverse Christoffel function $k_{n}$, the computation of which requires an orthonormal basis. As a result, existing approaches for non-orthogonal bases $\Phi_{n}$ typically rely on costly discrete orthogonalization. We propose a new iterative algorithm, inspired by recent advances in approximate leverage score sampling, that avoids this bottleneck. Crucially, while the computational and memory cost of discrete orthogonalization scales with $\|k_{n}\|_{L^{\infty}(X)}$, the cost of our approach increases only logarithmically in $\|k_{n}\|_{L^{\infty}(X)}$. In addition, we account for finite-precision effects by considering a numerical variant of the Christoffel function, ensuring that the algorithm relies only on computable quantities. Alongside a convergence proof, we present extensive numerical experiments demonstrating the efficiency and robustness of the proposed method.
@@ 1 Introduction
We consider the problem of approximating a function $f\in L^{2}_{\rho}(X)$, defined on $X\subseteq\mathbb{R}^{d}$, using weighted least squares fitting in the span of a given set of functions $\Phi_{n}=\{\phi_{1},\dots,\phi_{n}\}$
@@
where the set of sample points $\{x_{k}\}_{k=1}^{m}$ and the weights $\{w_{k}\}_{k=1}^{m}$ can be chosen freely. The objective is to attain near-best approximation errors while using a small number of sample points.
A common strategy is to use sample points that are drawn i.i.d. from some sampling distribution, allowing for an analysis using tools from random matrix theory. This led to the sampling distribution $\mu_{n}$, first proposed for (total degree) polynomial spaces [hampton2015coherence] and later generalized [cohen2017optimal], which achieves a near-optimal approximation error in probability using only $m=\mathcal{O}(n\log n)$ samples. This is close to the optimal case of interpolation, which corresponds to m=nm=n [^1] In the introduction, we assume for simplicity of exposition that $\dim\left(\operatorname{span}\left(\{\phi_{i}\}_{i=1}^{n}\right)\right)=n$, i.e., that $\Phi_{n}$ forms a basis. This assumption is not required and is relaxed later.. The sampling distribution $\mu_{n}$ is defined by
@@
is known as the inverse of the Christoffel function [nevai1986geza] and, hence, this sampling strategy is often referred to as Christoffel sampling [adcock2024optimal].
@@
In earlier work [dolbeault2022optimal, adcock2020near, migliorati2021multivariate], this issue was addressed by discretely orthogonalizing the basis $\Phi_{n}$ on a dense grid. More specifically, we can equivalently characterize the inverse Christoffel function $k_{n}$ using the non-orthogonal basis $\Phi_{n}$ by
@@
where the points $\{t_{k}\}_{k=1}^{\ell}$ are sampled i.i.d. from the measure $\rho$. Following [dolbeault2022optimal, adcock2024optimal], one needs $\ell=\mathcal{O}(\|k_{n}\|_{L^{\infty}(X)}\log(n))$ for $G_{\ell}$ to approximate $G$ well with high probability. Although this is typically an offline cost—computed once for every $\Phi_{n}$ and independently of the function to be approximated—it can still be prohibitively expensive, especially when the Christoffel function $k_{n}$ is sharply peaked. Notably, this is precisely the regime where Christoffel sampling offers significant advantages over standard least squares fitting using samples drawn from $\rho$. When $\ell$ is too small, the efficiency of the resulting Christoffel sampling procedure is limited [trunschke2024optimal].
We adopt a different strategy inspired by [cohen2015uniform]; Section 1.2 explains how this discrete method relates to our continuous setting. We aim to construct a function $u$ that satisfies $u(x)\geq k_{n}(x),\forall x\in X$ while $\|u\|_{L^{1}_{\rho}(X)}\leq Cn$ for some small constant $C\geq 1$. When such a function is known, it suffices to draw $m=\mathcal{O}(\|u\|_{L^{1}_{\rho}(X)}\log(n))=\mathcal{O}(n\log(n))$ samples from the sampling distribution $\mu_{u}$ defined by
@@
in order to obtain near-optimal weighted least squares approximations in probability. We construct such a function $u$ efficiently using an iterative method in which, at each step, we sample points from the current $u$ and thereafter use these samples to update and improve $u$.
In particular, we show that $\|u\|_{L^{1}_{\rho}(X)}$ decreases by a constant factor in each iteration until it is close to $\|k_{n}\|_{L^{1}_{\rho}(X)}=n$, using only $\mathcal{O}(n\log(n))$ samples per step. Starting from the initialization $u(x)=\|k_{n}\|_{L^{\infty}(X)}$, it follows that the total number of samples required to obtain an adequate $u$ is of the order $\mathcal{O}(\log(\|k_{n}\|_{L^{\infty}(X)})n\log(n))$. Therefore, in contrast to the method described above, the computational complexity of this approach depends only logarithmically on $\|k_{n}\|_{L^{\infty}(X)}$, rather than linearly. Inspired by [cohen2015uniform, Algorithm 3], the algorithm is termed refinement-based Christoffel sampling (RCS).
@@ 1.1 Effects of finite-precision arithmetic
Another issue related to least squares fitting with non-orthogonal bases is the influence of numerical rounding errors. Even though every basis is linearly independent from an analytical perspective, it can be indistinguishable from a linearly dependent set in finite-precision arithmetic. Numerically computed least squares fits using such a basis behave either explicitly or implicitly as $\ell^{2}$-regularized approximations with a regularization parameter $\epsilon\propto\epsilon_{\text{mach}}\sqrt{\|G\|_{2}}$ due to rounding errors [adcock2019frames, adcock2020frames, herremans2025sampling]. Whereas this numerical regularization has a negligible influence if the basis is (nearly) orthogonal, it becomes significant when $\kappa(G)\geq 1/\epsilon_{\text{mach}}^{2}$, in which case $\Phi_{n}$ is called numerically redundant [herremans2025sampling].
Using similar tools from random matrix theory, it was found in [herremans2025sampling] that the sampling distribution $\mu_{n}^{\epsilon}$ defined by
@@ 1.2 Connection to the discrete setting
Christoffel sampling is widely recognized as being closely related to leverage score sampling, a well-established technique in numerical linear algebra [drineas2006sampling, adcock2024optimal]. Given an $m\times n$ matrix $A$ with $m>n$, the $j$-th statistical leverage score $\ell_{j}$ is defined by
@@
where $U$ denotes the $m\times n$ matrix containing the left singular vectors of the matrix $A$ [drineas2012fast]. Leverage scores identify which rows of a matrix are most important and are typically used to subsample a tall-and-skinny matrix $A$ before solving a least squares problem $Ax\approx b$. Other applications include, but are not limited to, low-rank approximation, matrix completion and kernel-based machine learning [drineas2016randnla].
@@
The numerical Christoffel function $k_{n}^{\epsilon}$ can similarly be seen as a continuous analogue of ridge leverage scores [herremans2025sampling], which arise in the context of $\ell^{2}$-regularized least squares problems [alaoui2015fast].
Challenges similar to those encountered in Christoffel sampling also appear in the context of leverage scores. In particular, computing leverage scores requires the full singular value decomposition of $A$, which is as computationally expensive as solving the original least squares problem. Consequently, there has been substantial research on fast computation—or rather, approximation—of leverage scores. One approach introduced in [cohen2015uniform, cohen2017input] consists of computing a tight upper bound to the (rigde) leverage scores via an efficient iterative scheme. This approach can be extended to the continuous setting and forms the foundation of this paper.
@@ 1.3 Other related work
Other methods have been proposed for efficiently computing the Christoffel function from a non-orthogonal basis. For example, in  [dolbeault2022optimal, §4] a multilevel strategy is introduced that approximates the inverse Christoffel function through successive refinements based on nested subspaces. However, this approach relies on rather strong assumptions about these subspaces and does not account for the effects of finite-precision arithmetic. Another iterative method is presented in [trunschke2024optimal], which requires the strong assumption $\lambda_{\min}(G)\gg\epsilon_{\text{mach}}$—where $\lambda_{\min}$ indicates the smallest nonzero eigenvalue—for computability.
Moreover, we note that optimizations have been studied to improve upon the Christoffel sampling scheme. These strategies reduce the number of required samples from $\mathcal{O}(n\log(n))$ to $\mathcal{O}(n)$, typically by leaving the i.i.d. setting. For instance, in [bartel2023constructive] a subsampling scheme is introduced that is able to compress a good sampling set, while the algorithm in [dolbeault2024randomized] is designed to directly sample from a distribution that incorporates both the Christoffel function and effective resistance. Implementing the latter can be done via rejection sampling with draws from the Christoffel function, as mentioned in [dolbeault2024randomized, §5], which makes our approach directly applicable in this setting as well.
The link between leverage scores and the Christoffel function is not new; a more extensive comparison can be found in [meyer2023near, §2]. Furthermore, in the discrete setting many more algorithms have been proposed to efficiently approximate leverage scores, such as [drineas2012fast] which uses random sketching to speed up the computations. Viewing the object $\begin{bmatrix}\phi_{1}&\phi_{2}&\dots&\phi_{n}\end{bmatrix}$ as the limit of a tall-and-skinny matrix was popularized in the Matlab package chebfun, where it is referred to as a quasimatrix [driscoll2014chebfun].
@@ 1.4 Overview of the paper
The remainder of the paper is organized as follows. In Section 2, the symbols and notation used throughout the paper are summarized. In Section 3, we introduce the proposed iterative algorithm, outline its main objective and provide an initial discussion of its computational performance. Section 4 presents a theoretical analysis of the method, including a proof of its convergence properties. Section 5 is devoted to extensive numerical experiments that illustrate the efficiency and robustness of our approach. Finally, Section 6 lists some conclusions and open problems.
@@ 2 Symbols and notation
in (1). $\Phi_{n}=\{\phi_{1},\dots,\phi_{n}\}\subset L^{2}_{\rho}(X)$ is a given set of non-orthogonal functions, which may even be linearly dependent. The vector of evaluations of $\Phi_{n}$ at a point $x\in X$ is denoted by $\phi(x)=\begin{bmatrix}\phi_{1}(x)&\dots&\phi_{n}(x)\end{bmatrix}^{\top}\in\mathbb{C}^{n}$. The continuous Gram matrix associated with $\Phi_{n}$ is an $n\times n$ matrix $G$ defined by
@@ 3.1 Main objective
The goal is to identify a small number of samples $\{x_{k}\}_{k=1}^{m}$ and weights $\{w_{k}\}_{k=1}^{m}$ so that the discrete least squares fit to a function $f$ in a given set $\Phi_{n}$ has near-optimal approximation error. Our approach to obtain such samples is to compute a function $u$ that satisfies
@@
closely approximates the continuous Gram matrix $G$ using only $m=\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))$ samples. This is formalized in Lemma 3.1, which generalizes [herremans2025sampling, Theorem 10].
@@
Let $\Phi_{n}=\{\phi_{i}\}_{i=1}^{n}\subset L^{2}_{\rho}(X)$ be a set of functions, where $X\subseteq\mathbb{R}^{d}$, $\epsilon\geq 0$, and $u$ be an integrable function satisfying
@@
This follows from a straightforward generalization of the proof of [herremans2025sampling, Theorem 10] using a generic $\Delta$ instead of $\Delta=1/2$, and from
@@
Throughout this paper we will assume that $k_{n}^{\epsilon}(x)>0,\forall x\in X$. This is a standard assumption in Christoffel sampling schemes and is satisfied if at every point $x\in X$ at least one function $\phi_{i}$ is nonzero. This condition is mild and, for example, holds whenever the constant function is included. An alternative is to slightly modify the Christoffel density, as explained in [adcock2024optimal, §6.1].
Using [herremans2025sampling, Theorem 2], this immediately implies that numerically computed least squares fits are stable and accurate.
@@
where the term $\epsilon^{2}/\lambda_{1}(G)$ is negligble when $\Phi_{n}$ is close to orthogonal. From here, it follows that the discrete least squares error is near-optimal using $\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))=\mathcal{O}(n\log(n))$ samples, by arguments analogous to those used in standard results such as [cohen2017optimal, Theorem 2.1].
@@ 3.2 Iterative scheme
The core strategy for constructing a function $u$ that satisfies (4) is an adaptation of the (ridge) leverage score sampling algorithm introduced in [cohen2015uniform, cohen2017input] to the continuous setting. Suppose we have a candidate function $u$ that upper bounds $k_{n}^{\epsilon}$ but with $\|u\|_{L^{1}_{\rho}(X)}\gg n^{\epsilon}$. Drawing fewer than $\mathcal{O}(\|u\|_{L^{1}_{\rho}(X)}\log(n^{\epsilon}))$ samples from the associated measure $\mu_{u}$ (3) is generally insufficient for accurate least squares fitting. However, these samples can still be used to improve uu—that is, to construct a new function $u^{(new)}$ that still upper bounds $k_{n}^{\epsilon}$ but has a reduced $L_{\rho}^{1}(X)$-norm. This procedure is applied iteratively, producing successively tighter upper bounds to $k_{n}^{\epsilon}$ until the requirement (4) is met.
@@
with high probability. Thus, for given $C_{1}$ and $\Delta$, choosing $C_{2}$ appropriately ensures a geometric decrease in $L_{\rho}^{1}(X)$-norm of $u$, using only $\mathcal{O}(n^{\epsilon}\log(n^{\epsilon}))$ samples in each iteration. Note that we do not have to compute $m$ or $\alpha$; we simply draw $\lceil\alpha m+1\rceil$ samples in each iteration.
@@
$\Phi_{n}\subset L^{2}_{\rho}(X)$, (upper bounds on) $\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}$ and $n^{\epsilon}$, parameters $C_{1},C_{2}>0$ and $0<\Delta<1$